two golfers, smith and jones, are attempting to qualify for a golf championship. it is estimated that the probability of jones qualifying is 0.8, and that probability of both smith and jones qualifying is 0.6. given that the probability of smith qualifying and the probability of jones qualifying are independent, find the probability that only one of them are qualify
Answer by Boreal(15235)

p(S) unknown
p(J)=0.8
p(J)*p(S)=0.6, because independent.
Therefore, p(S)=0.75
Probability that only one will qualify is P(J)*p(NS)=0.8*0.25=0.2+
p(NJ)(p(S)=0.2*0.75=0.15
They add to 0.35